# A Decade of Systems for Human Data Interaction

### v1

@@ 3.3 From Single Table to Multi-Table Analytics
Instead, data engineers predefine denormalized views that join relevant tables [denorm, denorm2], typically managed through so-called semantic layers. This incurs substantial storage overhead if the joins are fully materialized, or introduces latency if executed per interaction.
Although not applied to visualization systems, the theoretical foundation to support interactive analytics over joins is well established. Probabilistic graphical models are akin to count-based analytics over a join graph. As early as the 1980s, the machine learning literature developed calibrated junction trees [shafer_shenoy1988, pearl1988probabilistic] as a partial aggregation data structure to support fast, exact marginal computations. We developed the Calibrated Junction Hypertree [Huang2023LightweightMF] to extend these ideas to any semiring aggregate and designed a set of reuse techniques to support OLAP and join operations (add, remove, change joins) in interactive latencies.
This system serves as the basis for designing join-oriented interactions and interfaces. Supporting multi-table semantics at the system layer also revealed subtle correctness errors that current interfaces silently inherit. Aggregations over 1-N or N-M joins can unintentionally duplicate data in ways that corrupt statistics and lead to incorrect visualization contents [Huang2023AggregationCE, hyde2024measures]. For example, ad-attribution pipelines often join users to impressions, causing user-level metrics to be replicated across many impressions and leading to inflated aggregates. A principled solution requires the system to track record multiplicities and for the interface to surface or infer the user’s intended aggregation semantics, so that input entity contributes exactly once.
@@ 3.4 From Single Table to Database Visualization
While the previous subsection showed how to perform interactive analytics over multi-table schemas, it focused on the data layer: visualization design still assumes a single-table input. Nearly all graphical grammars—rooted in Bertin’s data–visual mappings and formalized in the grammar of graphics [Wilkinson2001nViZnA, Wickham2010ALG, stolte2002polaris, Satyanarayan2018VegaLiteAG, wilkinson2012grammar]—treat a visualization as a mapping from a table: tuples become marks, attributes become visual channels, and the schema determines the structure of the view.
@@ 3.4.1 Illustrative Example: Node-Link Layouts
This limitation affects both how multi-table visualizations are specified (as above) and which designs are even expressible. Designs such as parallel coordinates, Sankey diagrams, or space-filling layouts internally encode multi-table structure but do so outside any visualization grammar, and instead rely on custom code or ad-hoc conventions to preserve relational semantics. What is needed is a principled extension from “Table Visualization” to Database Visualization [dvlicdt, dvlarxiv].
@@ 4 What About AI?
Large language models (LLMs) are sequential neural networks pre-trained on web-scale text corpora [vaswani2017attention]. They have demonstrated impressive capabilities in language understanding, reasoning, domain knowledge, programming [yang2024sweagent], and data tasks [shankar2024docetl, tan2024large, demiralp2024making]. Abstractly, an LLM acts as a function $f(any)\to any$: any task that can be serialized as text can be used as input and interpreted as output. Yet, this apparent generality hides a deep uncertainty. User goals may be under-specified, prompt descriptions imprecise, and the model’s outputs are probabilistically sampled from large, unstructured distributions. As a result, guarantees of safety, correctness, and performance remain elusive.
Despite these limitations, LLMs are already being used to augment or replace parts of data analysis and visualization pipeline, including data modeling [kotelnikov2023tabddpm, zhang2023mixed], SQL generation [satriani2025logical], dashboard creation [dibia2023lida, wang2025data, chen2024viseval], exploratory analysis [dibia2023lida], and even product ideation [productstrategy]. This breadth naturally raises the question: if LLMs can seemingly do everything, what role remains for HDI systems research?
To explore this question through the lens of interfaces, consider the rapidly growing practice of vibe coding [vibecoding]: generating functional artifacts through natural-language interaction with an LLM or agent. Tools such as Cursor, Replit, Lovable, and Convex Chef have gained rapid adoption by lowering the barrier to implementing data workflows. An Anthropic blog post reports using Claude Code to accelerate infrastructure, product design, growth, and data science development [claudecode]. They state that “despite knowing ‘very little JavaScript and TypeScript,’ the team uses Claude Code to build entire React applications for visualizing RL model performance and training data. These examples suggest that LLMs can now synthesize non-trivial user interfaces for complex data tasks, dramatically reducing the friction of manual UI development.
However, these interfaces still depend on robust backend systems. As Section 2 illustrated, even achieving interactive latencies is technically demanding, let alone ensuring consistency, reliability, or safety. LLMs lack the performance models, hardware awareness, and correctness invariants needed to make systems-level trade-offs: storage layout, physical operator choices, concurrency, and cache reuse. Empirical studies confirm these gaps. LLM accuracy collapses as task complexity grows [lin2025zebralogic, illusion]; models fail to apply even simple compiler rewrites [fang2024llmpeep]; and they often reason shallowly about programs, with poor generalization [cummins2023llmcomp]. LLM-generated code can orchestrate components, but guarantees must come from the underlying systems.
@@ 6 Acknowledgements
We thank Jiaxiang Liu, Charlie Summers, Haonan Wang, Lampros Flokas, Fotis Psallidas, Thibault Sellam, Qianrui Zhang, Haoci Zhang, Yifan Wu, Larry Xu, Daniel Alabi, Ziyun Wei, Lana Ramjit, Jacob Fisher, Jeff Tao for their contributions to the research lab, ideas, and the cited projects. We are grateful to Remco Chang, Joe Hellerstein, Arnab Nandi, Yunhai Wang, Subrata Mitra, Ravi Netravali, Aditya Parameswaran, Javad Ghaderi, Dan Rubenstein for collaborations and formative discussions over the years.

### v2

@@ 3.3 From Single Table to Multi-Table Analytics
Instead, data engineers predefine denormalized views that join relevant tables [44, 6], typically managed through so-called semantic layers. This incurs substantial storage overhead if the joins are fully materialized, or introduces latency if executed per interaction.
Although not applied to visualization systems, the theoretical foundation to support interactive analytics over joins is well established. Probabilistic graphical models are akin to count-based analytics over a join graph. As early as the 1980s, the machine learning literature developed calibrated junction trees [69, 57] as a partial aggregation data structure to support fast, exact marginal computations. We developed the Calibrated Junction Hypertree [37] to extend these ideas to any semiring aggregate and designed a set of reuse techniques to support OLAP and join operations (add, remove, change joins) in interactive latencies.
This system serves as the basis for designing join-oriented interactions and interfaces. Supporting multi-table semantics at the system layer also revealed subtle correctness errors that current interfaces silently inherit. Aggregations over 1-N or N-M joins can unintentionally duplicate data in ways that corrupt statistics and lead to incorrect visualization contents [35, 39]. For example, ad-attribution pipelines often join users to impressions, causing user-level metrics to be replicated across many impressions and leading to inflated aggregates. A principled solution requires the system to track record multiplicities and for the interface to surface or infer the user’s intended aggregation semantics, so that each input entity contributes exactly once.
@@ 3.4 From Single Table to Database Visualization
While the previous subsection showed how to perform interactive analytics over multi-table schemas, it focused on the data layer: visualization design still assumes a single-table input. Nearly all graphical grammars—rooted in Bertin’s data–visual mappings and formalized in the grammar of graphics [82, 81, 72, 68, 83]—treat a visualization as a mapping from a table: tuples become marks, attributes become visual channels, and the schema determines the structure of the view.
@@ 3.4.1 Illustrative Example: Node-Link Layouts
This limitation affects both how multi-table visualizations are specified (as above) and which designs are even expressible. Designs such as parallel coordinates, Sankey diagrams, or space-filling layouts internally encode multi-table structure but do so outside any visualization grammar, and instead rely on custom code or ad-hoc conventions to preserve relational semantics. What is needed is a principled extension from “Table Visualization” to Database Visualization [87, 85].
@@ 4 What About AI?
Large language models (LLMs) are sequential neural networks pre-trained on web-scale text corpora [76]. They have demonstrated impressive capabilities in language understanding, reasoning, domain knowledge, programming [88], and data tasks [70, 73, 19]. Abstractly, an LLM acts as a function $f(any)\to any$: any task that can be serialized as text can be used as input and interpreted as output. Yet, this apparent generality hides a deep uncertainty. User goals may be under-specified, prompt descriptions imprecise, and the model’s outputs are probabilistically sampled from large, unstructured distributions. As a result, guarantees of safety, correctness, and performance remain elusive.
Despite these limitations, LLMs are already being used to augment or replace parts of data analysis and visualization pipeline, including data modeling [45, 89], SQL generation [66], dashboard creation [20, 78, 12], exploratory analysis [20], and even product ideation [61]. This breadth naturally raises the question: if LLMs can seemingly do everything, what role remains for HDI systems research?
To explore this question through the lens of interfaces, consider the rapidly growing practice of vibe coding [43]: generating functional artifacts through natural-language interaction with an LLM or agent. Tools such as Cursor, Replit, Lovable, and Convex Chef have gained rapid adoption by lowering the barrier to implementing data workflows. An Anthropic blog post reports using Claude Code to accelerate infrastructure, product design, growth, and data science development [4]. They state that “despite knowing ‘very little JavaScript and TypeScript,’ the team uses Claude Code to build entire React applications for visualizing RL model performance and training data. These examples suggest that LLMs can now synthesize non-trivial user interfaces for complex data tasks, dramatically reducing the friction of manual UI development.
However, these interfaces still depend on robust backend systems. As Section 2 illustrated, even achieving interactive latencies is technically demanding, let alone ensuring consistency, reliability, or safety. LLMs lack the performance models, hardware awareness, and correctness invariants needed to make systems-level trade-offs: storage layout, physical operator choices, concurrency, and cache reuse. Empirical studies confirm these gaps. LLM accuracy collapses as task complexity grows [47, 71]; models fail to apply even simple compiler rewrites [23]; and they often reason shallowly about programs, with poor generalization [18]. LLM-generated code can orchestrate components, but guarantees must come from the underlying systems.
@@ 6 Acknowledgements
We thank Jiaxiang Liu, Charlie Summers, Haonan Wang, Lampros Flokas, Fotis Psallidas, Thibault Sellam, Qianrui Zhang, Haoci Zhang, Viraj Rai, Yifan Wu, Larry Xu, Daniel Alabi, Xiaolan Wang, Niranjan Kamat, Lilong Jiang, Gabriel Ryan, Ziyun Wei, Zhengjie Miao, Marianne Procopio, Ab Mosca, Lana Ramjit, Jacob Fisher, Zhaoning Kong, Jeff Tao for their contributions to the research lab, ideas, and the cited projects. We are grateful to Remco Chang, Joe Hellerstein, Arnab Nandi, Yunhai Wang, Subrata Mitra, Ravi Netravali, Jiannan Wang, Alexandra Meliou, Leilani Battle, Sam Madden, Carlos Scheidegger, Aditya Parameswaran, Javad Ghaderi, Dan Rubenstein for collaborations and formative discussions over the years.